### v2

@@ 2.2 Experimental design and estimation
$\mathcal{C}_{n}$ is measurable with respect to $\mathbf{A}$ and is such that the number of units in cluster $k$ is $n_{k}=\gamma_{k}\frac{n}{K_{n}}$ with $\max_{k}\gamma_{k}\leq\bar{\gamma}<\infty$, for some $\bar{\gamma}<\infty$. Assume that $D_{i}=\tilde{D}_{c(i)}$ almost surely, where $\tilde{D}_{c(i)}|\mathbf{A},\{Y_{i}(\mathbf{d})\}_{i\in\{1,\cdots,n\}\mathbf{d}\in\{0,1\}^{N}}\sim\mathrm{Bern}(0.5)$ are independent across clusters.
@@
Motivated by standard practice both in industrial applications and in field experiments with clusters (Baird et al., 2018), we consider estimators obtained by simple differences in means between treated and control clusters. Because $P(D_{i}=1)=1/2$, we construct a (biased) estimator of treatment effects as
@@
Note that $Y_{1}-\bar{\mu}_{1},\ldots,Y_{n}-\bar{\mu}_{n}$ typically have different ranges ex ante even under Assumption 2. To account for this, we can apply the more general result in Section 6.2. ∎
Remark 5 (Alternative estimators).
Alternative estimators studied in the literature are inverse probability weights estimators (e.g., Aronow and Samii, 2017; Ugander et al., 2013) (these estimators reweight by the probability that the treatment of all friends being either zero or one). Unless researchers impose additional restrictions on the exposure mapping, in the network context, these estimators can be subject to the instability of the propensity score because they reweight by the inverse probability that all friends of a given individual are either under treatment or control. An alternative are model-based estimators, which are subject to model misspecification. A third class of estimators are trimming estimators, which trim observations based on their position in the network. We show how our framework generalizes for such estimators in Section 6.3. ∎
@@
It is possible to consider designs where treatments are assigned with different cluster-level probabilities. Our main insights for the bias and variance characterization follow similarly to what we present for cluster experiments, with appropriate modifications presented in Section 6.4. ∎
Figure 1: Example of clustering design with a single network. The network is partitioned into three clusters. Elements in a given cluster are assigned the same color.
3 (When) should you cluster?
This section studies when a given clustering $\mathcal{C}_{n}$ should be chosen over a simple Bernoulli design where treatments are randomized independently across individuals. This is often a central question in experiments (e.g. Holtz et al., 2020). To study this question, we first characterize the bias and variance of the estimator $\hat{\tau}_{n}(\mathcal{C}_{n})$.
We denote a Bernoulli design as a special cluster design with clusters $\mathcal{C}_{\mathrm{B},n}=\{c_{k}=\{k\},k\in\{1,\cdots,n\}\}$. Denote $\mathbb{E}_{\mu}[\cdot]$ the expectation for given potential outcome functions $\mu$ (conditional on $\mathbf{A}$). The components $\mathcal{O}(\cdot),o(\cdot)$ in the following lemmas and theorems hold uniformly over all cluster designs with $K_{n}$ many clusters that satisfy Assumption 4.
Our goal is to design experiments that minimize the worst-case variance, while controlling the worst-case bias. For a given clustering $\mathcal{C}_{n}$, the dual of the optimization problem is:
$\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)=\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}+\lambda\sup_{\mu\in\mathcal{M}}\Big{(}\tau_{n,\mu}-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2},$ (5)
where the supremum is intended over $(\mu_{i})_{i=1}^{n}$, and $\mathcal{M}$ is the product space of potential outcome functions as in Assumption 2. The parameter $\lambda$ is user-specific and denotes the relative importance weight assigned to the worst-case bias. For $\lambda=1$, Theorem 3.4 below shows that $\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)$ coincides in the limit with the worst-case mean-squared error.
@@ 3.1 Worst-case bias, variance, and MSE
As a first step, we characterize the worst-case bias of the estimator $\hat{\tau}_{n}(\mathcal{C}_{n})$. Let
@@
Let Assumptions 1, 2, 4 hold. Then
$\sup_{\mu\in\mathcal{M}}\Big{|}\tau_{n,\mu}-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{|}=\bar{\phi}_{n}b_{n}(\mathcal{C}_{n}).$
@@
See Appendix C.2. ∎
Lemma 3.1 shows that the worst-case bias can be expressed as the number of friends of $i$ in a different cluster from $i$, appropriately reweighted by the overall number of friends of $i$. The size of the worst-case bias also depends on the magnitude of spillover effects $\bar{\phi}_{n}$. Lemma 3.1 provides a formal justification to notions of between clusters connectedness based on the worst-case bias of the treatment effect estimator.
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}\times\Big{[}\bar{\psi}+o(1)\Big{]}.$
@@
See Appendix C.5. ∎
Lemma 3.2 characterizes the worst-case variance as the number of clusters $K_{n}$ grows (at an arbitrary rate). Such bounds depend on the variance in cluster sizes, $\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2},$ times a positive constant $\bar{\psi}$. In Section 7 we provide practical recommendations for the choice of $\bar{\psi}$. By leveraging the local asymptotic framework in Assumption 3, a key insight is that the variance is (asymptotically) driven by the within-cluster correlations instead of cross-cluster connections. This intuition formalizes the idea that a larger variation in clusters’ size increases the variance of the estimator. To our knowledge, these characterizations of the worst-case bias and variance are novel to the literature on spillover effects.
Combining Lemma 3.1 and Lemma 3.2, we obtain the expression of $\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)$.
@@
Suppose that Assumptions 1, 2, 3, 4 hold. Then as $K_{n}\rightarrow\infty$
$\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)=\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)(1+o(1)),$
where
@@
Theorem 3.3 is a direct corollary of Lemmas 3.1, 3.2. ∎
An alternative quality measure of the estimator is the worst-case mean square error (MSE), defined as
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}[(\hat{\tau}_{n}(\mathcal{C}_{n})-\tau_{n,\mu})^{2}]$ (9)
Since the MSE is obtained by adding up the square of bias and variance, it is evidence that the worst-case MSE is upper bounded by $\mathcal{B}_{n}(\mathcal{C}_{n},1)$. Here, we can also show that they coincide as $K_{n}\rightarrow\infty$. [^13] The worst-case MSE has been considered as the criterion in other settings for design of randomized experiments (Wu, 1981).
Theorem 3.4.
Suppose that Assumptions 1, 2, 3, 4 hold. Then as $K_{n}\rightarrow\infty$,
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}[(\hat{\tau}_{n}(\mathcal{C}_{n})-\tau_{n,\mu})^{2}]=\mathcal{B}_{n}^{*}(\mathcal{C}_{n},1)(1+o(1)).$
Proof.
See Appendix C.6. ∎
Remark 7 (Unobserved $\mathbf{A}$).
Suppose that $\mathbf{A}$ is unobserved or partially observed, and researchers have a prior over $\mathbf{A}$. In this case, the worst-case variance does not change, the worst-case bias is replaced by $\bar{\phi}_{n}\mathbb{E}_{\mathbf{A}}[b_{n}(\mathcal{C}_{n})]$, and $b_{n}^{2}(\mathcal{C}_{n})$ in $\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)$ is replaced by $\mathbb{E}_{\mathbf{A}}[b_{n}^{2}(\mathcal{C}_{n})]$, where the expectation may depend on partial network information (Breza et al., 2020). ∎
3.2 Comparison with a Bernoulli design
Theorem 3.5.
$\liminf_{n\rightarrow\infty}\frac{\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)}{\mathcal{B}_{n}(\mathcal{C}_{\mathrm{B},n},\lambda)}>1\quad\text{ if }\sqrt{K_{n}}\bar{\phi}_{n}\rightarrow 0\text{ and }K_{n}/n\leq\kappa,$ (10)
$\limsup_{n\rightarrow\infty}\frac{\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)}{\mathcal{B}_{n}(\mathcal{C}_{\mathrm{B},n},\lambda)}<1\quad\text{ if }\sqrt{K_{n}}\bar{\phi}_{n}\rightarrow\infty,\text{ and }b_{n}(\mathcal{C}_{n})\leq\delta,$
where $b_{n}(\mathcal{C}_{n})$ is defined in (6).
Proof.
See Appendix C.7. ∎
Theorem 3.5 compares a Bernoulli design and a cluster design. Theorem 3.5 states that we should not run a cluster experiment if the size of the spillover effects $\bar{\phi}_{n}$ goes to zero at an order faster than $1/\sqrt{K_{n}}$, where $K_{n}$ denotes the number of clusters, and the number of clusters is sufficiently smaller than the sample size (at an arbitrary rate). We should instead run a cluster design if spillover effects are larger in magnitude than $1/\sqrt{K_{n}}$. The assumption that $b_{n}(\mathcal{C}_{n})\leq\delta$ is equivalent to assuming that $\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}|\mathcal{N}_{i}\bigcap\{j:c(j)\neq c(i)\}|\leq\delta$, for some $\delta\in[0,1)$, i.e., the clustering decreases the bias of the Bernoulli design (since $b_{n}(\mathcal{C}_{\mathrm{B},n})=1$).
Table 1: Implications of Theorem 3.5 and examples of rates.
Description Formulation Implication: Run Examples of Rates
small ##\## of clusters
very small ##\## of clusters
large ##\## of clusters
Table 2: Minimum spillover effects’ size that justifies running a cluster experiment instead of a Bernoulli design in the presence of equally sized clusters $(\gamma_{k}=1$ for all $k\in\{1,\cdots,K_{n}\})$. The bias denotes $\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}|j\in\mathcal{N}_{i}:c(i)\neq c(j)|\in[0,1]$. The table on the left-hand side reports the value of $\bar{\phi}_{n}$ for different values of the bias and number of clusters, fixing $\bar{\psi}=4$ (e.g., valid when outcomes are bounded between $[-1,1]$). In the figure on the right-hand side, the y-axis reports in log-scale the minimum size of the spillover effects $\bar{\phi}_{n}/\sqrt{\bar{\psi}}$ to run a cluster experiment as in Theorem 3.6. The x-axis reports the number of clusters. Different colors report different values of the bias.
@@ 3.2 Comparison with a Bernoulli design
To complement the properties as a function of rates of convergence, in the following theorem, we provide an explicit rule of thumb for a cluster design. It will be useful to define
$\xi_{n}=(\lambda\bar{\phi}_{n}^{2}/\bar{\psi})^{-1},$ (11)
denoting the ratio of the largest deviation $\bar{\psi}$ over the magnitude of the spillovers squared times the weight $\lambda$ on the bias. Our algorithm and explicit recommendations will require researchers to specify $\xi_{n}$ as a tuning parameter (see Section 7 for recommendations for $\xi_{n}$).
Theorem 3.6 (Rule of thumb).
Suppose that Assumptions 1, 2, 3, 4 hold. Let $\underline{\gamma}=\frac{1}{K_{n}}\sum_{k=1}^{K_{n}}\gamma_{k}^{2}$, where $\gamma_{k}=n_{k}K_{n}/n$. Then for some $N>0,n>N$
$\frac{\mathcal{B}_{n}^{*}(\mathcal{C}_{\mathrm{B},n},\lambda)}{\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)}\geq 1\quad\text{ if }\quad\xi_{n}\leq\frac{K_{n}(1-b_{n}(\mathcal{C}_{n})^{2})}{\underline{\gamma}}.$
Proof of Theorem 3.6.
See Appendix C.8. ∎
Theorem 3.6 provides a rule of thumb for choosing a cluster design with given clusters $\mathcal{C}_{n}$ over a Bernoulli design $\mathcal{C}_{\mathrm{B},n}$, using the approximate objective $\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)$. The right-hand side depends on three observables: (i) $\underline{\gamma}$, (ii) the expected bias of the clustering method (as a function of $\mathbf{A}$), and (iii) the number of clusters $K_{n}$. The left-hand side depends on $\xi_{n}$. For $\lambda=1$, known $\bar{\psi}$, the rule of thumb provides the smallest spillover effects that would guarantee that the cluster design dominates the Bernoulli design. When clusters are equally sized, the minimum size of the spillover effects to run a cluster experiment is $\bar{\phi}_{n}\geq\sqrt{\frac{\bar{\psi}}{(1-b_{n}(\mathcal{C}_{n})^{2})K_{n}}}$, where $b_{n}(\mathcal{C}_{n})$ is defined in (6).
For example, assume (i) equally sized clusters, (ii) the bias of the clustering is at most $50\%$ as a conservative upper bound, and (iii) outcomes are bounded between zero and one (in which case $\bar{\psi}\leq 4$). In this setting, researchers should run a cluster experiment when $\bar{\phi}_{n}\sqrt{K_{n}}$ is larger than 2.3. Figure 2 illustrates the exact threshold for the minimum size of spillovers that justifies conducting a cluster experiment. In our empirical application in Section 5 we find $\bar{\phi}_{n}\geq 0.003$ would motivate conducting a cluster experiment. [^14] This magnitude is constistent with typical magnitudes of treatment and spillover effects obtained from information campaigns on online platforms (Karrer et al., 2021; Athey et al., 2023).
4 Choosing the clusters
In this section, we turn to designing the optimal clustering. Theorem 3.3 characterizes the objective function. After simple re-arrangement, Theorem 3.3 provides a simple-to-compute metric for ranking (a few) given clusters (with $\xi_{n}$ as in Equation 11)
$\frac{\xi_{n}}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}+\Big{(}\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}\Big{|}j\in\mathcal{N}_{i}:c(i)\neq c(j)\Big{|}\Big{)}^{2}.$ (12)
When practitioners must choose between a few different clusters, we recommend drawing the frontier in Equation (12) as a function of the bias $\Big{(}\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}\Big{|}j\in\mathcal{N}_{i}:c(i)\neq c(j)\Big{|}\Big{)}^{2}$ and the variance $\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}$, with multiplier over the variance $\xi_{n}=(\lambda\bar{\phi}_{n}^{2}/\bar{\psi})^{-1}$. Clusters that perform reasonably well compared to other clusters for a large set of values of $\xi_{n}$ should be preferred for implementation.
4.1 Causal clustering: algorithm
Equation (12) is computationally difficult to optimize over a large space of clusters. In particular, the task of estimating the best clustering over a large class is challenging because the number of connections between different clusters enters non-linearly in Equation (12). First, we consider an objective as a function of the absolute (instead of squared) bias.
Corollary 4.1.
Suppose that the conditions in Theorem 3.3 hold. Then as $K_{n}\rightarrow\infty$
$\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)\leq\lambda\bar{\phi}_{n}^{2}\left(\frac{\xi_{n}}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}\Big{[}1+o(1)\Big{]}+\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}\Big{|}\mathcal{N}_{i}\bigcap\Big{\{}j:c(i)\neq c(j)\Big{\}}\Big{|}\right).$ (13)
Corollary 4.1 shows that Equation (13) with $\xi_{n}$ as in Equation (11), as a function of the absolute instead of squared bias, is a surrogate (upper bound) loss of the objective function.
We interpret the objective in Equation (12) as minimizing the worst-case variance under constraints on the squared worst-case bias (whose dual depends on the multiplier $\lambda$). Whenever we are interested in the frontier that trade-offs the bias and variance over different values of $\xi_{n}$, Equation (12) or (13) have the same dual representation for (different) values of the multipliers $\xi_{n}$ in each equation. Corollary 4.1 shows that we can choose the same $\xi_{n}$ to obtain an upper bound on the original objective.
Following Corollary 4.1, we optimize over $(K_{n},c(\cdot))$ as a function of a constant $\xi_{n}$,
$\frac{\xi_{n}}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}+\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}\Big{|}j\in\mathcal{N}_{i}:c(i)\neq c(j)\Big{|}.$ (14)

### v3

@@ 2.2 Experimental design and estimation
$\mathcal{C}_{n}$ is a deterministic function of $\mathbf{A}$ and is such that the number of units in cluster $k$ is $n_{k}=\gamma_{k}\frac{n}{K_{n}}$ with $\max_{k}\gamma_{k}\leq\bar{\gamma}<\infty$, for some $\bar{\gamma}<\infty$. Assume that $D_{i}=\tilde{D}_{c(i)}$ almost surely, where for $k\in\{1,\cdots,K_{n}\}$, $\tilde{D}_{k},|\mathbf{A},\{Y_{i}(\mathbf{d})\}_{i\in\{1,\cdots,n\}\mathbf{d}\in\{0,1\}^{N}}\sim\mathrm{Bern}(0.5)$ and $\tilde{D}_{1},\cdots,\tilde{D}_{K_{n}}|\mathbf{A},\{Y_{i}(\mathbf{d})\}_{i\in\{1,\cdots,n\}\mathbf{d}\in\{0,1\}^{N}}$ are independent across clusters.
@@
Motivated by standard practice both in industrial applications and in field experiments with clusters, we consider estimators obtained by simple differences in means between treated and control clusters (which are those used in most of applications on online platforms, Karrer et al., 2021); see Remark 3 for comparison with other estimators studied in the literature. Because $P(D_{i}=1)=1/2$, we construct a (biased) estimator of treatment effects as [^15] Here, for expositional convenience, we make explicit the dependence of $\hat{\tau}_{n}$ with $\mathcal{C}_{n}$, where the clustering $\mathcal{C}_{n}$ only characterizes the distribution of treatment assignments $D_{i}$ but not the structure of the estimator.
@@
Remark 3 (Alternative estimators).
Alternative estimators studied in the literature are inverse probability weights estimators (e.g., Aronow and Samii, 2017; Ugander et al., 2013). These estimators reweight by the probability that the treatment of all friends being either zero or one. Unless researchers impose additional restrictions on the exposure mapping, in the network context, these estimators are highly sensitive to instability of the propensity score because the propensity score scale to zero exponentially in the maximum degree. [^16] Even if the maximum degree grows at a slow rate, such as $\mathcal{N}_{n,\max}\propto\log(n),\bar{\phi}_{n}\propto 1/\log^{2}(n)$, so that effects are local to zero but non negligible for inference the propensity score converges to zero at rate $1/n$. An alternative are model-based estimators, which are subject to model misspecification. A third class of estimators are trimming estimators, which trim observations based on their position in the network and also optimize over the choice of the estimators and designs. We study this class of estimators and designs in Appendix B.3. ∎
@@
Note that $Y_{1}-\bar{\mu}_{1},\ldots,Y_{n}-\bar{\mu}_{n}$ typically have different ranges ex ante even under Assumption 2. To account for this, we can apply the more general result in Appendix B.2. ∎
@@
It is possible to consider designs where treatments are assigned with different cluster-level probabilities. Our main insights for the bias and variance characterization follow similarly to what we present for cluster experiments, with appropriate modifications presented in Appendix B.4. ∎
3 Bias, Variance and MSE of a given clustering
As a necessary first step, this section focuses on studying the bias, variance, and mean-squared error of a given clustering $\mathcal{C}_{n}$, deferring the question of choosing an optimal clustering to the next section. Denote $\mathbb{E}_{\mu}[\cdot]$ the expectation for given potential outcome functions $\mu$ (conditional on $\mathbf{A}$). The components $\mathcal{O}(\cdot),o(\cdot)$ in the following lemmas and theorems hold uniformly over all cluster designs with $K_{n}$ many clusters that satisfy Assumption 4.
Our goal is to design experiments that minimize a weighted combination of the variance and squared bias, worst-case over the class of potential outcomes.
Our objective reads as follows
$\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)=\sup_{\mu\in\mathcal{M}}\left\{\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}+\lambda\Big{(}\tau_{n,\mu}-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\right\},$ (5)
where the supremum is intended over $(\mu_{i})_{i=1}^{n}$, and $\mathcal{M}$ is the product space of potential outcome functions as in Assumption 2. The user may specify $\lambda$.
For $\lambda=1$, $\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)$ coincides in the limit with the worst-case mean-squared error. It is also possible to study optimal clustering to minimize the length of confidence intervals for the overall effects, which we defer to Section 6.1.
@@ 3.1 Worst-case bias, variance, and MSE
As a preliminary step in the characterization of the objective function, we first characterize the worst-case bias and variance. Let
@@
Let Assumptions 1, 2, 4 hold. Then $\sup_{\mu\in\mathcal{M}}\Big{|}\tau_{n,\mu}-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{|}=\bar{\phi}_{n}b_{n}(\mathcal{C}_{n}).$
@@
See Appendix D.2. ∎
Lemma 3.1 shows that the worst-case bias can be expressed as the number of friends of $i$ in a different cluster from $i$, appropriately reweighted by the overall number of friends of $i$. The size of the worst-case bias also depends on the magnitude of spillover effects $\bar{\phi}_{n}$. Lemma 3.1 provides a formal justification to notions of between clusters connectedness based on the worst-case bias of the treatment effect estimator. In the proof, the worst-case bias is attained as the condition in Equation (16) holds with equality (Appendix B.2 studies settings where effects are heterogeneous across individuals). The expression for the bias connects to measures of cluster purity such as those in Goel and Salganik (2009), Toulis and Kao (2013), motivating clustering purity based on the class of models in Assumption 2.
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}\times\Big{[}\bar{\psi}+\mathcal{O}(\varepsilon)\Big{]}.$
@@
See Appendix D.5. ∎
Lemma 3.2 characterizes the worst-case variance as the number of clusters $K_{n}$ grows (at an arbitrary rate). Such bounds depend on the variance in cluster sizes, $\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2},$ times a positive constant $\bar{\psi}$. By leveraging the local asymptotic framework in Assumption 3, a key insight is that the variance is (asymptotically) driven by the within-cluster correlations instead of cross-cluster connections, up-to a $\varepsilon$-approximation. To our knowledge, this characterization of the worst-case variance is novel to the literature.
@@
Suppose that Assumptions 1, 2, 3, 4 hold. Then as $K_{n}\rightarrow\infty$ $\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)=\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)(1+\mathcal{O}(\varepsilon)),$ where
@@
Proof.
See Appendix D.6. ∎
Remark 6 (Choice and sharper bound for $\bar{\psi}$).
Because here we consider a finite population with fixed potential outcomes, the worst-case variance depends on $\bar{\psi}$, a constant proportional to the largest value of the realized outcomes. In Appendix B.1, we provide similar calculations in settings where potential outcomes are random functions, in which case we can interpret $\bar{\psi}$ as a sharper bound on the moments of potential outcomes. ∎
Remark 7 (Unobserved $\mathbf{A}$).
Suppose that $\mathbf{A}$ is unobserved or partially observed, and researchers have a prior over $\mathbf{A}$. In this case, the worst-case variance does not change, the worst-case bias is replaced by $\bar{\phi}_{n}\mathbb{E}_{\mathbf{A}}[b_{n}(\mathcal{C}_{n})]$, and $b_{n}^{2}(\mathcal{C}_{n})$ in $\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)$ is replaced by $\mathbb{E}_{\mathbf{A}}[b_{n}^{2}(\mathcal{C}_{n})]$, where the expectation may depend on partial network information (Breza et al., 2020). In this case, the sparsity restrictions in Assumption 1 can be replaced by assumptions on the expected squared degree over the prior – see Equation (38) in the proof of Lemma D.2. ∎
3.2 Comparison between a given clustering and a Bernoulli design
Before discussing the optimal choice of the cluster, we explore a simpler question, that is central in many randomized trials with spillovers. Suppose we are given a clustering $\mathcal{C}_{n}$, but we have the option to run a Bernoulli design where treatments are assigned independently between individuals. When should be prefer the (more complex) cluster design over a Bernoulli design? Denote a Bernoulli design as $\mathcal{C}_{\mathrm{B},n}=\{c_{k}=\{k\},k\in\{1,\cdots,n\}\}$.
Theorem 3.4 (Rule of thumb).
Suppose that Assumptions 1, 2, 4 hold. Let $\underline{\gamma}=\frac{1}{K_{n}}\sum_{k=1}^{K_{n}}\gamma_{k}^{2}$, where $\gamma_{k}=n_{k}K_{n}/n$. Then
$\frac{\mathcal{B}_{n}^{*}(\mathcal{C}_{\mathrm{B},n},\lambda)}{\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)}\geq 1\quad\text{ if }\quad(\lambda\bar{\phi}_{n}^{2}/\bar{\psi})^{-1}\leq\frac{K_{n}(1-b_{n}(\mathcal{C}_{n})^{2})}{\underline{\gamma}}.$
@@ 3.2 Comparison between a given clustering and a Bernoulli design
See Appendix D.7. ∎
Theorem 3.4 provides an explicit rule for choosing a cluster design with given clusters $\mathcal{C}_{n}$ over a Bernoulli design $\mathcal{C}_{\mathrm{B},n}$, using the approximate objective $\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)$. The right-hand side depends on three observables: (i) $\underline{\gamma}$, (ii) the expected bias of the clustering method (as a function of $\mathbf{A}$), and (iii) the number of clusters $K_{n}$. The left-hand side depends on $\bar{\phi}_{n}^{2}/\bar{\psi}$. For $\lambda=1$, known $\bar{\psi}$, the rule of thumb provides the smallest spillover effects that would guarantee that the cluster design dominates the Bernoulli design. Figure 1 illustrates the exact threshold for the spillovers that justifies a cluster experiment. In our empirical application in Section 5 we find $\bar{\phi}_{n}\geq 0.003$ would motivate a cluster experiment.
Table 1: Minimum spillover effects’ size that justifies running a cluster experiment instead of a Bernoulli design in the presence of equally sized clusters $(\gamma_{k}=1$ for all $k\in\{1,\cdots,K_{n}\})$, and where the clustering $\mathcal{C}_{n}$ is given. The bias denotes $\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}|j\in\mathcal{N}_{i}:c(i)\neq c(j)|\in[0,1]$. The table on the left-hand side reports the value of $\bar{\phi}_{n}$ for different values of the bias and number of clusters, fixing $\bar{\psi}=4$ (e.g., valid when outcomes are bounded between $[-1,1]$). In the figure on the right-hand side, the y-axis reports in log-scale the minimum size of the spillover effects $\bar{\phi}_{n}/\sqrt{\bar{\psi}}$ to run a cluster experiment as in Theorem 3.4. The x-axis reports the number of clusters.
@@
4 Optimization over the choice of the clustering
In this section, we turn to designing the optimal clustering.
As a first observation, we note that Theorem 3.3 implies that any clustering solving the original objective in Equation (5) must be on the Pareto frontier for the worst-case bias and variance as derived in Lemmas 3.1 and 3.2, respectively. Therefore, we first focus on drawing the frontier between the variance and bias. The optimization problem to draw the frontier between the worst-case bias and variance can be equivalently expressed in terms of the absolute instead of quadratic bias, so that we study
$L_{c}(\xi)=\frac{\xi}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}+\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}\Big{|}j\in\mathcal{N}_{i}:c(i)\neq c(j)\Big{|}.$ (9)
We are interested in minimizing $L_{c}(\xi)$ over a grid of values for $\xi$. This effectively will allow us to draw a Pareto frontier between the worst-case variance and absolute bias. Any clustering solving the original objective in Equation (5) must minimize $L_{c}(\xi)$ for some value of $\xi$. Once we obtain the Pareto frontier, computing the clustering that minimizes our original objective in Equation (5), becomes a simple task, as it only requires comparing the objective in Equation (5) within the frontier.